OpenStudy (sparklestaraa):

Marco and his two younger sisters would like to purchase a silver charm bracelet for their mother’s birthday. They went to the mall and found what they were looking for at Store A. In Store A, the bracelet without charms costs $85 and each charm costs $15.

OpenStudy (sparklestaraa):

A. Use function notation that models the total price of the bracelet and how that price is based on the number of charms. Explain the reasoning behind your equation. B. What would be a reasonable domain for this function based on this scenario? Explain why this is a proper domain. C. If Marco and his sisters have saved $250, make a graph to show all of the charms they could purchase with a bracelet if they had saved between $85 and $250. D. Marco found five charms that he likes. Using your function, determine the cost of the bracelet he would make. E. As they walked down the hall of the mall, Marco saw this sales flyer for Store B on a bench.

OpenStudy (sparklestaraa):

@misty1212

OpenStudy (sparklestaraa):

Marco is wondering from which store he should buy the items so that he spends the least amount of money. Which store should Marco use if he wanted the five charms and silver bracelet all together? What advice can you tell Marco to help him shop wisely depending on the number of charms he wants to buy? Justify how you reasoned your answers.

OpenStudy (misty1212):

HI!

OpenStudy (sparklestaraa):

Haiii :D

OpenStudy (misty1212):

if you buy say \(x\) charms then you have to pay \(\$15\) each for a total cost of \[C(x)=85+15x\]

OpenStudy (sparklestaraa):

thats A?

OpenStudy (misty1212):

lol yeah that is A

OpenStudy (misty1212):

the domain would be the number of charms you buy you can buy 0, 1, 2, 3, 4,5, ... charms so i would say whole number from zero to say 20 or so?

OpenStudy (misty1212):

and that is B

OpenStudy (sparklestaraa):

for b just say that it would be a whole number from zero to 20 or more?

OpenStudy (misty1212):

If Marco and his sisters have saved $250, make a graph to show all of the charms they could purchase with a bracelet if they had saved between $85 and $250. i don't know how to make a graph, but the most charms they could buy you would solve \[85+15x=250\] and get \[15x=165\\ x=11\]so between zero charms and eleven charms

OpenStudy (misty1212):

you decide how many charms you can fit on one bracelet for the domain

OpenStudy (misty1212):

i just made up the 20

OpenStudy (misty1212):

Marco found five charms that he likes. Using your function, determine the cost of the bracelet he would make.\[C(5)=85+15\times 5\] compute that number

OpenStudy (sparklestaraa):

how would i make this graph?

OpenStudy (sparklestaraa):

oh oh u said u dont know xd

OpenStudy (misty1212):

no idea i guess you could graph the line \[y=85+15x\] http://www.wolframalpha.com/input/?i=y%3D85%2B15x

OpenStudy (sparklestaraa):

x=11? for the first

OpenStudy (sparklestaraa):

omg im so behind xd ur telling me it and I'm just blindly answering like u never told me xd

OpenStudy (misty1212):

but don't make it look like that because it can't be negative make it look like this

OpenStudy (misty1212):

go slow, read what i wrote, it is all there be careful

OpenStudy (misty1212):

i have answered A, B, C, and D after you read them let me know, make sure you understand them, then we can do E if you like

OpenStudy (sparklestaraa):

oh so dont make it look like that graph?

OpenStudy (misty1212):

no because the domain is not all real numbers, they have to be positive. only what it looks like in the first quadrant i can't seem to use the draw tool to show you

OpenStudy (misty1212):

nope, not working

OpenStudy (sparklestaraa):

would d be 20?

OpenStudy (sparklestaraa):

ok back up to D I'm still confused on that which thing did u write that is D?

OpenStudy (misty1212):

D i wrote here Marco found five charms that he likes. Using your function, determine the cost of the bracelet he would make. \[C(5)=85+15×5\] compute that number

OpenStudy (sparklestaraa):

oh and i got 20 for c ?

OpenStudy (misty1212):

\[85+15\times 5\]

OpenStudy (sparklestaraa):

160 :D

OpenStudy (misty1212):

ok but C is not 20 i wrote it above, question and answer take a look

OpenStudy (sparklestaraa):

15x=165 x= 11 that one?

OpenStudy (misty1212):

yes that one

OpenStudy (sparklestaraa):

oh ur saying they can buy 11 charms @misty1212

OpenStudy (misty1212):

yes zero to 11 because that is what you get if you solve \[85+15x=250\]for \(x\)

OpenStudy (sparklestaraa):

so ok 11 charms i would make the graph starting at 85 on the y axis and 0-11 on the x axis and plot dots on each like 0-11 or would i use a graph?

OpenStudy (misty1212):

sure why not?

OpenStudy (sparklestaraa):

https://www.desmos.com/calculator if i type in the equation do i get the graph?

OpenStudy (sparklestaraa):

@misty1212

OpenStudy (misty1212):

this is going to give you a lousy picture the scale is wrong and you only want the graph in quadrant i

OpenStudy (sparklestaraa):

oh ok let me make a pic and see if its right

OpenStudy (misty1212):

k

OpenStudy (sparklestaraa):

no this won't work i was going to do a line graph but no @misty1212

OpenStudy (misty1212):

yeah i know i tried too just move on

OpenStudy (sparklestaraa):

So d is 160 ryt? @misty1212

OpenStudy (misty1212):

YUP

OpenStudy (sparklestaraa):

:D yay ok so lets move on to E

OpenStudy (misty1212):

the first one we had the equation \[85+15x\] the second one would be \[100+12x\]and you want to know which one is cheaper it will depend on the number of charms you buy if you buy a lot the second one will be if you buy fewer, the first one will be cheaper to find the number at which it changes set them equal and solve for x

OpenStudy (sparklestaraa):

how do i set them as equal do i put it like 85+15x 100+15x ?

OpenStudy (misty1212):

\[85+15x=100+12x\] yes

OpenStudy (sparklestaraa):

ok so I'm still very confused

OpenStudy (sparklestaraa):

like Which store should Marco use if he wanted the five charms and silver bracelet all together? What advice can you tell Marco to help him shop wisely depending on the number of charms he wants to buy? Justify how you reasoned your answers. thats what it asks and I'm still confused on how i would tell him how to shop wisely etc.

OpenStudy (sparklestaraa):

@misty1212

OpenStudy (misty1212):

did you solve the equation ?

OpenStudy (sparklestaraa):

x=5?

OpenStudy (misty1212):

yeas good work

OpenStudy (misty1212):

now lets go slow ok?

OpenStudy (sparklestaraa):

kk

OpenStudy (sparklestaraa):

C is the graph @DanJS

OpenStudy (misty1212):

a) if they buy 5 charms, they pay the same amount of money, they pay $160 we computed that already in an earlier problem b) if they by less than 5, the one that cost \(85+15x\) will be cheaper, because the the price with no charms is cheaper, and it take up until 5 to make the price the same as the other one c) if they buy more than 5, the one that cost \(100+12x\) is cheaper, because that one cost less per charm

OpenStudy (danjs):

why u need me?

OpenStudy (sparklestaraa):

graph for C

OpenStudy (danjs):

you figured the cost function C as a function of number of charms x C(x) = 15x + 85

OpenStudy (danjs):

If Marco and his sisters have saved $250, make a graph to show all of the charms they could purchase with a bracelet if they had saved between $85 and $250.

OpenStudy (danjs):

\[$85 \le c(x) \le $250\]

OpenStudy (sparklestaraa):

equation for the graph?

OpenStudy (danjs):

What is the flyer for? that is part e) right

OpenStudy (sparklestaraa):

yes

OpenStudy (danjs):

ok, ill graph it , one sec

OpenStudy (sparklestaraa):

thanks :D

OpenStudy (danjs):

the red line is the cost function y-axis is the cost x-axis is # charms for the cost to be between 85 and 250, the number of charms has to be between 0 and 11

OpenStudy (danjs):

that look alright to you?

OpenStudy (sparklestaraa):

Yes :D omg thanks soo much :D